OpenStudy (anonymous):

2(k + 4) - 3k < 14 (but the < is underlined)

OpenStudy (anonymous):

you mean \[2(k+4) - 3k \leq 14\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

\[2k + 8 - 3k \leq 14\]\[-k \leq 6\]\[k \geq -6\]

OpenStudy (anonymous):

and when do you have to change (reverse) the signs?

OpenStudy (lilg132):

\[2k + 8 - 3k \le 14\] \[-k \le 14 - 8\] \[-k \le 6\] \[k \ge -6\]

OpenStudy (anonymous):

you mean how \[-k \leq 6\]became \[k \geq -6\]? There's one way to think about it

OpenStudy (anonymous):

yes that :)

OpenStudy (lilg132):

when the variable is a minus then you just flip the sides over to change the variable to a positive

OpenStudy (anonymous):

oh okay! :)

OpenStudy (anonymous):

\[-k \leq 6\]to get rid of that negative sign on k, you can think of switching the left-hand-side and the right-hand-side by adding k to both sides, and then subtracting 6 from both sides \[0 \leq 6 + k\]\[-6 \leq k\]then it becomes clear that \[-6 \leq k\]is the same thing as \[k \geq -6\]

OpenStudy (anonymous):

thanks! :)

OpenStudy (lilg132):

your welcome :)